The distance between a town and factory is 30Km. A man started to walk From the factory to the town at 6:30 am. While a cyclist left the town for the factory at 6:50 am riding at a speed of 18Km/hr , the man met the cyclist after walking 6Km . Find at what time they met..

Grade:8

1 Answers

Arun
6 years ago
Let the cyclist start from town (A) & man started from factory (B)


Distance travelled by a cyclist to reach point C= 30-6=24 km

Speed of cyclist= 18 km/h

Time taken by the cyclist to reach point C= Distance/speed


Time taken by the cyclist to reach point= 24/18= 4/3 h = (4/3)×60= 80 min= 1h 20 min 

The required time at which they will meet= 6:50 + 1:20 = 7h 70 min= 8 :10 a.m

Hence, The required time at which they will meet= 8:10 a.m
 
Regards
Arun (askIITians forum expert)